Write a system of linear equations to represent the situation. Then, answer the question using (a) a table, (b) a graph, and (c) algebra.
2. You buy 16 candles. Large candles cost $7 each and small candles cost $3 each. You spend $76. How many large candles x and small candles y did you buy?
Number of candles: Number of large candles plus number of small candles is 16.
Cost of candles: $7 times number of large candles plus $3 times number of small candles is $76

Answer by ikleyn(52781)

.
You are on the right way:

L + S = 16      (1)   (large candles and small candles)
7L + 3S = 76    (2)   (dollars).

Express L = 16 - S from (1) and substitute it into (2), replacing L.
You will get a single equation for S:

7*(16-S) + 2S = 76.


Can you complete on your own from this point ?

This method is called the "Substitution Method".